Week of 060417

Open Actions from last week: Review Easter weekend problems

Chair: James

On Call: James/Maarten

Monday

Log: Vacation

Actions:

Discussion:

Tuesday:

Log: Weekend ok - 1.6GB/s av.

New Actions:

  • Bernd Panzer requests 6 hours continuation of disk-disk while he adds in writing to the sc4 disk pool. - DONE
  • Switch to tape tomorrow -DONE
Discussion:

Wednesday:

Log: disc-disc finished. 1GB rate to tape in same pool shown.

New Actions:

  • switchover of LFC aliases to new hosts -DONE
Discussion:

Thursday

Log:nothing to report

Actions:

Discussion:

Friday

Log: nothing to report

Actions:

Discussion:

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2006-04-21 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback